the sides of a parallelogram are in ratio 2:3 find the adjacent sides of parallelogram , if the perimeter of the parallelogram is 20 cm.
Answer by flame8855(424)

y/x = 2/3
x+y=10
2/3x +x =10
5/3x=10
x =6
y=4